HUBBLE SPACE TELESCOPE DAILY REPORT #2273 PERIOD COVERED: 0000Z (UTC) 12/17/98 - 0000Z (UTC) 12/18/98 Daily Status Report as of 352/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy VCC1413 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Four Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.5 Completed Eleven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.6 Completed FGS 7492 (The Distances to Dwarf Novae, and the Calibration of the Technique of Infrared Spectroscopic Parallax) The Fine Guidance Sensors were used to make astrometric observations of the dwarf nova SS-Cyg and several other reference stars. The distances to dwarf novae are poorly known. The distance to the object can be measured using a technique of infrared spectroscopic parallax. The observations were completed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Six Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.9 Completed Two Sets of NIC/2 7828 (Detection of the Infrared Jet in the R-Aquarii Binary System) The Near Infrared Camera (Camera Two) was used to make observations of the variable binary star R-Aqrarii to detect its infrared jet. The nearby {~200 pc} R-Aqrarii symbiotic system is a well-known prototype for non-relativistic astrophysical jets. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 6751 (Is the Central Star of K 648 in M15 a Close Binary?) The WF/PC-2 was used to make observations of the central star of the planetary nebula K648 which is located in globular cluster M-15 to determine if this star is a binary star. Pre-repair HST images reveal that the nebula is strongly elliptical in shape, and shows blobs on either side of the central star. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Two Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Three Sets of WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxies 3C326.1, 3C119.0 and 3C137.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.13 Completed STIS/CCD/MA1 8067 (Improved STIS Sensitivity Measurement for E140H) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of G191B2B. This program is an improved sensitivity measurement for the E140H grating. A wavelength calibration observations was also completed. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed Two Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.16 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.18 Completed WF/PC-2 7274 (Secular Changes in the Jet of M-87) The WF/PC-2 was used to make observations of the secular changes in the jet coming off of the active galactic nuclei of galaxy M-87. Past observations with the FOC have shown changes as small as 0.03", and provide a unique opportunity to study proper motions and structural changes of the jet. An internal flat calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.19 Completed WF/PC-2 6587 (Black Holes and Cores of Early-Type Galaxies) The WF/PC-2 was used to make observations of early-type galaxy IC3370 in order to study its nuclear core region and its possible black hole. The campaign to find black holes in the centers of galaxies using HST has so far been highly successful. The observations were completed as planned, and no problems were reported. 1.20 Completed STIS/CCD 8066 (IR Standards: HZ43 Repeat) The Space Telescope Imaging Spectrograph (CCD) was used to make infrared observations of HZ43. Only a few calibrations measurements exist, more data are needed, especially in G230LB, which has no data at all. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 13 Successful : 13 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 31 Successful : 30 The FHST update at 352/0633Z failed. 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared six times. FHST performance data was collected twice per an open operations request. As documented in HSTAR 6819, an incorrect return link frequency was automatically scheduled for an SSR DG1 TDRSS support. This resulted in a 18-minute, 1-second loss of recoverable engineering data. The PACOR-II facility reported, via a TTR, three PSN jumps at 352/0234Z. At 352/045622Z, a STIS MCE-2 reset occurred outside a SAA contour while the low voltage was on. HSTAR 6820 was written. ROP NS-2 was executed at 352/0513Z to clear the STIS flight software error counter. Recovery was accomplished via SMS commanding at 352/051338Z, the time of MAMA-2 high voltage on commanding. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH